OpenStudy (anonymous):

If ƒ(x )= 2x^2 + 3, then ƒ(x - 1)= 2x^2 + 4x + 5. True False

OpenStudy (anonymous):

@manumben

OpenStudy (anonymous):

plug in 2(x-1)^2 + 3, and simplify.

OpenStudy (anonymous):

Idk how to do these

OpenStudy (anonymous):

2((x-1)(x-1))+3

OpenStudy (anonymous):

..?

OpenStudy (anonymous):

are you familiar with the FOIL concept?

OpenStudy (anonymous):

(x-1)(x-1) = x^2 - 2x +1

OpenStudy (anonymous):

2x^2-4x+2+3 = 2x^2-4x+5

OpenStudy (anonymous):

is 2x^2-4x+5 = 2x^2 +4x + 5

OpenStudy (anonymous):

Thanks